. In Harrod Domar model, suppose the savings rate is 20% and capital output is 4. The

population rate of growth is 4%. What is going to be the growth path of this economy?

Explain



1
Expert's answer
2021-03-09T15:19:14-0500

If the savings rate is 20% and capital output is 4, then the growth path of this economy is G = s/C = 20/4 = 5%.